OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2 .

OpenStudy (anonymous):

What are the equations?

OpenStudy (anonymous):

they are y − 6 =1/2 (x + 3) y =1/2 x − 6 y + 3 =1/2 (x − 6) x − 2y = 12

hero (hero):

Hint: use y = mx + b to figure out the line. Since you're given the slope, m = 1/2 and the point (x,y) = (6,-3), substitute those into the formula to get -3 = 1/2(6) + b Then solve for b to get -3 = 3 + b -3 - 3 = b -6 = b So b = -6 Now re-write the line using only the slope and the y-intercept in the form y = mx + b: y = 1/2x - 6

OpenStudy (anonymous):

thank you